OpenStudy (danikasosa):

If f(x) = 4x – 5 and g(x) = 3 – 2x, find f(x) + g(x). A. 2x – 2 B. 6x – 2 C. 2x – 8 D. 6x – 8

OpenStudy (will.h):

Alone got this ;)

alones (alones):

Will got this (;

OpenStudy (will.h):

Never am here to watch you killing it ;)

OpenStudy (yournightmare):

a

alones (alones):

I believe it would be A, since you're addding them.

OpenStudy (will.h):

Told you Alone Got this

alones (alones):

Told you Will would supervise you

OpenStudy (will.h):

Lol. Enough courtesy they'll think of at in bad intention lool

OpenStudy (kels2153):

\[(4x-5)+(3-2x)=4x-2x-5+3\] Now, you simplify. \[4x-2x=2x\] and \[-5+3=-2\] Therefore the answer is: \[2x-2\]